OpenStudy (anonymous):

factor then use zero product principle 5x^3+3x^2=125x+75

OpenStudy (anonymous):

x^2(5x+3)=75(5x+3) ==> x^2=75 ==> x=5sqrt(5) or x=-5sqrt(5)

OpenStudy (anonymous):

x=5sqrt(3) or x=-5sqrt(3)

OpenStudy (anonymous):

take: 5x^3+3x^2-125x-75=0 (5x+3)(x^2-25)=0 5x+3=0===>x=-3/5 x^2-25=0===>(x+5)(x-5)===>x-5 and x=-5

OpenStudy (anonymous):

thats what i did thanks!! :D

OpenStudy (anonymous):

np